# Enhancing Vision-Language Compositional Understanding with Multimodal Synthetic Data

### v1

@@ 3.3 Training with Real and Synthetic Samples
where $m_{0}$ is a base margin, $\beta<0$ is a cutoff threshold, $\gamma$ is a scaling factor. The adaptive margin is designed with the following rationale: when $d<\beta$, the negative similarity difference, which is large in absolute value, suggests the presence of incorrect or mislabeled samples, hindering model training. In this case the margin is set to $d$, resulting in zero loss in Eq. (5). For $\beta\leq d\leq m_{0}$, the margin is scaled, with smaller differences receiving larger margins to emphasize learning from harder samples. When $d>m_{0}$, the samples are already well-separated with the margin $m_{0}$, so the margin is capped at $m_{0}$, leading to zero loss. We illustrate the adaptive margins and the loss for different values of $d$ in Figure 2 (b).
The loss for all images is given by $L_{mar}^{I}=\frac{1}{3n}\sum_{i}L_{mar,i}^{I}$. Similarly, we compute $L_{mar}^{T}$ for all captions, and the total adaptive margin loss is $L_{mar}=L_{mar}^{I}+L_{mar}^{T}$. Finally, the overall training loss is a weighted combination of the contrastive loss and the adaptive margin loss, with a weighting factor $\lambda$:
@@
Table 1: Comparison of accuracy (%) between SVD-GT and baselines on four benchmarks. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Source # real img # real cap # syn img # syn cap
@@ 4.1 Datasets
Evaluation. We evaluate our model on four vision-language compositional understanding benchmarks: (1) ARO [103], which consists of 23,937 cases for relation understanding and 28,748 for attribute understanding. We exclude the subsets for order understanding, as they contain significant nonsensical and non-fluent artifacts [32]. (2) VL-CheckList [110], a large-scale benchmark with over 100,000 samples, evaluates compositionality across subsets of objects, attributes, and relationships, which are further divided into various fine-grained categories. (3) SugarCrepe [32] includes 7,000 test cases across seven subsets. In the above three benchmarks, each test case containing one image, one positive caption, and one negative caption. (4) SugarCrepe++ [20] includes 4,757 test samples across five subsets, where each test case consists of one image, two positive captions, and one negative caption. All benchmarks involve classifying captions as positive or negative for the given images. We report the average accuracy across all subsets of each benchmark and include the accuracy for each subset in Appendix Sec. A3.
@@ 4.2 Implementation Details
Training Setups. We use the AdamW optimizer [55] with a cosine annealing learning rate schedule [54] to adjust the learning rates. Training is conducted on two Tesla V100 GPUs, with a batch size of 128 sample groups for ViT-B/32 and 16 for ViT-L/14. The base learning rate is set to 0.01 for a total batch size of 256, with adjustments made according to the Linear Scaling Rule [25] based on the actual batch size. Training is performed for 3,000 steps for ViT-B/32 and 15,000 steps for ViT-L/14, corresponding to fewer than 5 epochs in previous studies [19, 103, 106]. More details about data synthesis hyper-parameters selections are included in Appendix Sec. A2.
Table 2: Ablated performance (%) of SVD-GT. “SynCap” refers to synthetic captions, “SynImg” refers to synthetic images, “FeatInj” denotes image feature injection, and “CompSet” indicates the comparison sets used in Eq. (5).
@@ 4.3 Main Results
We compare SVD-GT with several baseline methods: (1) methods that utilize synthetic samples for training [48, 64, 76, 103, 80, 106, 105, 66], (2) methods that distill knowledge from pretrained models [5, 111, 77, 12], (3) methods that incorporate knowledge from scene graphs [82, 19], and (4) methods that leverage detailed image captions [11, 44]. Both SVD-GT and the baseline methods use ViT-B/32 as the base model. The results are presented in Table 1.
We observe that SVD-GT achieves the best performance on VL-CheckList, SugarCrepe, and SugarCrepe++ compared to baselines trained on the same data source, namely COCO. Specifically, SVD-GT surpasses the strongest baseline by 1.4% on VL-CheckList and 2.5% on SugarCrepe, underscoring its effectiveness in enhancing compositional understanding. However, SVD-GT performs worse on ARO than some baselines. We hypothesize that this may be due to nonsensical or grammatically incorrect artifacts in the ARO test samples [32], which could favor methods that utilize rule-based synthetic training samples (e.g., [82]) containing similar artifacts. Notably, even with only 50% of the training data, SVD-GT outperforms most baseline methods (see Table 3 for SVD-GT performance with reduced training data). Compared to baselines using additional data sources, SVD-GT surpasses or matches their performance. For example, SVD-GT outperforms methods in [19, 12] despite their use of more training data. SVD-GT performs comparably to CE-CLIP+ [106] on VL-CheckList and SugarCrepe, despite CE-CLIP+ leveraging CC3M as an additional data source and utilizing significantly more samples.
@@ 4.4 Ablation Study and Analysis
We perform ablation studies to assess the impact of each component and design choice in SVD-GT. The results, presented in Table 2, show the overall performance as the average accuracy across the four benchmarks, as well as the individual accuracy for each benchmark.
@@
How do synthetic captions and images influence compositional understanding? To analyze the impact of synthetic captions and images, we compare Variant #1 (using only real samples) with Variant #2 (using real samples and synthetic captions) and #3 (using real samples and synthetic images). Variant #2 shows a significant improvement over Variant #1, increasing the average accuracy from 69.33% to 73.99% for ViT-B/32. However, Variant #3 obtains only a marginal improvements of about 0.5%. These results suggest that synthetic captions substantially enhance compositional understanding, while synthetic images alone provide limited benefit, which aligns with findings in [65]. We hypothesize that generative artifacts in synthetic images are more noticeable than in captions and affect the effective learning of nuanced distinctions of VLMs, thereby reducing their effectiveness. When both synthetic captions and images are used (Variant #4), performance improves further, indicating a synergistic effect of combining both modalities.
@@
How do image feature injection and AdaIN contribute to compositional understanding? To analyze the impact of image feature injection and AdaIN, we compare Variant #4 (without image feature injection or AdaIN) with Variant #5 (with image feature injection only) and #6 (with AdaIN only). Using ViT-B/32, we observe that Variant #5 outperforms Variant #4 across all four benchmarks, with an average accuracy increase of 0.8%. While Variant #6 shows a smaller improvement of 0.4%, it outperforms Variant #5 on ARO and SugarCrepe. These results suggest that image feature injection is more effective for improving compositionality than AdaIN, although the two methods have some complementary effects. Through the improvements obtained using ViT-L/14 are a little bit different, both techniques improve performance over Variant #4. Combining both methods in Variant #7 leads to further improvements, with a more than 1% increase in average accuracy for ViT-B/32 and a 0.65% increase for ViT-L/14 over Variant #4. This highlights the importance of reducing unintended changes in synthetic images, as such changes could cause the model to rely on them for distinguishing positive and negative samples rather than the semantic differences specified in the corresponding captions. Examples of synthetic images in Figure A2 in Appendix show how image feature injection helps mitigate these unintended changes.
@@
What are the contributions of margin loss, and which adaptive margin strategy yields the best performance? To evaluate the impact of adaptive margin loss and its different strategies, we compare three variants with distinct margin approaches: (1) Fixed (Variant #8): a fixed margin without adaptive adjustments; (2) Adaptive (Variant #9): the margin is calculated according to Eq. (6); and (3) Adaptive Inversed (Variant #10): a larger margin is applied to samples with higher similarity differences, representing an inverse version of Eq. (6) that prioritizes learning from easier samples. Comparing Variant #8 with Variant #7, we observe that Variant #8 improves by over 1% for ViT-B/32 and 0.6% for ViT-L/14 in average accuracy, suggesting that margin loss effectively aids in distinguishing between positive and negative samples, thereby enhancing compositional understanding. Between Variants #8, #9, and #10, we find that Variant #9 outperforms Variant #8 across all four benchmarks, with an average accuracy improvement of over 0.3% for both ViT-B/32 and ViT-L/14. In contrast, Variant #10 results in a performance decline on three benchmarks and no improvement in average accuracy. These findings highlight the superior effectiveness of adaptive margins.
@@
Table 3: Performance (%) of SVD-GT using different subsets of training samples. “Neg” represents synthetic negative samples, “Pos” represents synthetic positive samples, and “Prop” represents the proportion of training samples.
Variant Neg Pos Prop   ARO VL-CheckList SugarCrepe SugarCrepe++ Average
@@
Do synthetic positive samples enhance compositional understanding? To assess the impact of synthetic positive samples, we compare the performance of training with only synthetic negative samples versus using both synthetic negative and positive samples. Based on Variant #7, we train ViT-B/32 with these two combinations of training data and present the results in Table 3. We observe a nearly 3% decrease in average accuracy when using only synthetic negative samples, compared to using both synthetic negative and positive samples. This decrease is particularly noticeable on SugarCrepe++, ARO, and VL-CheckList. These results underscore the importance of incorporating synthetic positive samples in training, as they provide essential information about variations that maintain semantic consistency, which is critical for compositional understanding.
How do model size and training data size influence compositional understanding? To evaluate the impact of model size on compositional understanding, we compare the performance of ViT-B/32 and ViT-L/14 in Table 2. We find that ViT-L/14 only provides a modest improvement of about 0.5% of average accuracy over ViT-B/32. This suggests that increasing model size does not significantly enhance compositional understanding. To investigate the effect of training data size, we train ViT-B/32 using 20% and 50% of randomly sampled training data and report the results in Table 3. We observe that performance improves as the proportion of training data increases, indicating that a larger training set helps the model better capture compositional knowledge. However, the performance gain diminishes when the training data size increases from 50% to 100%. Perhaps generating or selecting high-quality data is more effective than simply increasing the data size.
Additional experiments on hyperparameter analysis can be found in Appendix Sec. A3.
@@ 5 Conclusion
In this paper, we introduce SVD-GT to enhance the compositional understanding capabilities of VLMs by generating and training with synthetic data. To tackle two key challenges in using synthetic data—namely, the difficulty of generating accurate variations and the inconsistency in cross-modal alignment quality—SVD-GT integrates image feature injection into a T2I model to improve the quality of synthetic variations and introduces an adaptive margin loss to account for varying levels of cross-modal alignment for effectively learning nuanced distinctions. Experiments on four visual-language compositional understanding benchmarks demonstrate the effectiveness of SVD-GT.
@@
  • Section A1 presents additional details about SVD-GT.
@@
A1 Details of SVD-GT
The prompts used as input to the LLM for generating positive and negative captions are presented in Figure A1.
@@
Data Synthesis. For caption generation, we utilize the Llama-2-Chat 13B model [^1] https://huggingface.co/meta-llama/Llama-2-13b-chat, with the temperature set to 0.9, top-k set to 100, and top-p set to 0.9 for sampling. For image generation, we use the LCM model [^2] https://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 [57]. The pretrained CLIP ViT-L/14 [71] is used as the image feature extractor for injecting image features. We perform 8 inference steps with LCM to generate each image.
Hyperparameter Selection. First, we use only real training samples to select $\tau$ and $b$. The optimal values are determined by searching for the ones that minimize the training loss during the first training step, aiming to preserve the output distribution from the pretrained model. After searching, we set $\tau=0.01$ and $b=-30.0$. Next, we select the base learning rate, weight decay, and LoRA adapter rank based on performance on the COCO-2014 validation set, training exclusively on real samples. According to the performance on the validation set, these hyperparameter are set to a base learning rate of 0.01, weight decay of 0.5, and LoRA adapter rank of 16. Then, we construct a validation set composed of the CIFAR-10 [46] test set and a randomly selected 5% of samples from ARO-Attribute and ARO-Relation, to balance the performance on coarse-grained and fine-grained tasks. Using this validation set, we train the model on both real and synthetic samples and use the validation performance to determine the remaining hyperparameters: $m_{0}$, $\alpha$, $\beta$, $\gamma$ and $\lambda$. The effects of these hyperparameters are shown in Table A4.
@@
Table A1: Comparison of accuracy (%) between SVD-GT and baselines on ARO and VL-CheckList. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Table A2: Comparison of accuracy (%) between SVD-GT and baselines on SugarCrepe. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Table A3: Comparison of accuracy (%) between SVD-GT and baselines on SugarCrepe++. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Effects of image feature injection. In Figure A2, we present examples of synthetic images to illustrate how image feature injection helps mitigate unintended changes. By comparing synthetic images with and without feature injection, we observe that the injection helps generate images with shot types similar to those of the real image. For example, in the first row, the real image shows a wide shot of a girl, while the synthetic image without feature injection shows a close-up shot, despite aligning with the caption. With image feature injection, the synthetic image displays a wide shot, similar to the real image. A similar effect is observed in the second row. Additionally, image feature injection helps generate backgrounds that resemble the real image. In the third row, the real image features an indoor background, while the synthetic image without feature injection displays a outdoor street scene, which differs significantly. With image feature injection, the single-colored background makes the synthetic image more similar to the real one. These examples demonstrate that image feature injection can reduce some unintended variations not reflected in the caption, making it useful for training VLMs.
Figure A2: Examples of synthetic samples without and with image feature injection. Image feature injection primarily influences shot type (close-up vs. wide shots in the 1st and 2nd rows) and background scenes (indoor vs. outdoor in the 3rd row), making the synthetic images more closely resemble real images in these aspects.
Effect of hyperparameters. Table A4 presents the performance of SVD-GT with different hyperparameter settings. For $\lambda$, we observe that $\lambda=0.01$ achieves the highest average validation accuracy, leading us to select it for subsequent experiments. Similarly, for $\alpha$, the best performance is obtained with $\alpha=10.0$, which is used in other experiments. When evaluating different values of $m_{0}$, we find that $m_{0}=0.005$ yields the best results. Finally, we examine various combinations of $\beta$ and $\gamma$ and observe that $\beta=-0.02$ and $\gamma=1.0$ provide the best validation performance. Thus, this combination is selected as the optimal hyperparameter setting.
@@
Table A4: Performance of SVD-GT with different hyperparameters. “ARO-Rel” refers to the ARO-Relation validation subset, and “ARO-Att” refers to the ARO-Attribute validation subset, both consisting of a randomly selected 5% of the full set, as described in Sec. A2.

### v2

@@ 3.3 Training with Real and Synthetic Samples
where $m_{0}$ is a base margin, $\beta<0$ is a cutoff threshold, $\gamma$ is a scaling factor. The adaptive margin is designed with the following rationale: when $d<\beta$, the purported positive pair is much less similar than the purported negative pair, suggesting the presence of incorrect or mislabeled samples, so we zero out the loss by setting the margin to $d$. For $\beta\leq d\leq m_{0}$, the margin is scaled, with smaller differences receiving larger margins to emphasize learning from harder samples. When $d>m_{0}$, the samples are already well-separated with the margin $m_{0}$, so the margin is capped at $m_{0}$, again leading to zero loss. We visualize both the adaptive margin and the corresponding loss as functions of $d$ in Figure 2 (b).
The loss for all images is given by $L_{mar}^{I}=\frac{1}{3n}\sum_{i}L_{mar,i}^{I}$. Analogously, we compute $L_{mar}^{T}$ for all captions. The total adaptive margin loss is $L_{mar}=L_{mar}^{I}+L_{mar}^{T}$. Finally, the overall training loss is a weighted combination of the contrastive loss and the adaptive margin loss, with weight $\lambda$:
@@
Table 1: Comparison of accuracy (%) between SPARCL and baselines on four benchmarks. “img.” represents images, “cap.” represents captions, “syn.” represents synthetic data.
@@
Source # real img. # real cap. # syn. img. # syn. cap.
@@ 4.1 Datasets
Evaluation. We evaluate our model on four vision-language compositional understanding benchmarks: (1) ARO [106], which consists of 23,937 cases for relation understanding and 28,748 for attribute understanding. We exclude the subsets for order understanding, as they contain significant nonsensical and non-fluent artifacts [32]. (2) VL-CheckList [113], a large-scale benchmark with over 100,000 samples, evaluates compositionality across subsets of objects, attributes, and relationships, which are further divided into various fine-grained categories. (3) SugarCrepe [32] includes 7,000 test cases across seven subsets. In the above three benchmarks, each test case containing one image, one positive caption, and one negative caption. (4) SugarCrepe++ [19] includes 4,757 test samples across five subsets, where each test case consists of one image, two positive captions, and one negative caption. All benchmarks involve classifying captions as positive or negative for the given images. We report the average accuracy across all subsets of each benchmark and include the accuracy for each subset in Appendix A3.
@@ 4.2 Implementation Details
Training Setups. We use the AdamW optimizer [59] with a cosine learning rate schedule [58]. Training is conducted on two Tesla V100 GPUs, with a batch size of 128 sample groups for ViT-B/32 and 16 for ViT-L/14. The base learning rate is set to 0.01 for a total batch size of 256, and scaled linearly [24] based on the actual batch size. Training is performed for 3,000 steps for ViT-B/32 and 15,000 steps for ViT-L/14, corresponding to fewer than 5 epochs in previous studies [18, 106, 109]. More details and hyperparameters settings are in Appendix A2.
Table 2: Ablated performance (%) of SPARCL. “SynCap” refers to synthetic captions, “SynImg” refers to synthetic images, “FeatInj” denotes image feature injection, and “CompSet” indicates the comparison sets used in Eq. (3).
@@ 4.3 Main Results
We compare SPARCL with several baseline methods: (1) methods that utilize synthetic samples for training [49, 68, 79, 106, 83, 109, 108, 70], (2) methods that distill knowledge from pretrained models [3, 114, 80, 11], (3) methods that incorporate knowledge from scene graphs [85, 18], and (4) methods that leverage detailed image captions [9, 45]. Both SPARCL and the baseline methods use ViT-B/32 as the base model. The results are presented in Table 1.
We observe that SPARCL achieves the best performance on VL-CheckList, SugarCrepe, and SugarCrepe++ compared to baselines trained on the same data source, namely COCO. Specifically, SPARCL surpasses the strongest baseline by 1.4% on VL-CheckList and 2.5% on SugarCrepe, underscoring its effectiveness in enhancing compositional understanding. However, SPARCL performs worse on ARO than some baselines. We hypothesize that this may be due to nonsensical or grammatically incorrect artifacts in the ARO test samples [32], which could favor methods that utilize rule-based synthetic training samples (e.g., [85]) containing similar artifacts. Notably, even with only 50% of the training data, SPARCL outperforms most baseline methods (see Table 3 for SPARCL performance with reduced training data). Compared to baselines using additional data sources, SPARCL surpasses or matches their performance. For example, SPARCL outperforms methods in [18, 11] despite their use of more training data. SPARCL performs comparably to CE-CLIP+ [109] on VL-CheckList and SugarCrepe, despite CE-CLIP+ leveraging CC3M as an additional data source and utilizing significantly more samples.
@@ 4.4 Ablation Study and Analysis
We perform ablation studies to assess the impact of each component and design choice in SPARCL. The results, presented in Table 2, show the overall performance.
@@
Synthetic captions vs. synthetic images. To analyze the impact of synthetic captions and images, we compare Variant #1 (using only real samples) with Variant #2 (using real samples and synthetic captions) and #3 (using real samples and synthetic images). Variant #2 shows a significant improvement over Variant #1, with the average accuracy improving from 69.33% to 73.99% for ViT-B/32. However, Variant #3 obtains only a marginal improvements of about 0.5%. These results suggest that synthetic captions substantially enhance compositional understanding, while synthetic images alone provide limited benefit, which aligns with findings in [69]. We hypothesize that generative artifacts in synthetic images are more pronounced than in captions, which negatively affects the effective learning of nuanced distinctions by VLMs. When both synthetic captions and images are used (Variant #4), performance improves further, indicating a synergistic effect.
@@
Image feature injection vs. AdaIN. To analyze the impact of image feature injection and AdaIN, we compare Variant #4 (without image feature injection or AdaIN) with Variant #5 (with image feature injection only) and #6 (with AdaIN only). Using ViT-B/32, we observe that Variant #5 outperforms Variant #4 across all four benchmarks, with an average gap of 0.8%. While Variant #6 shows a smaller improvement of 0.4%, it outperforms Variant #5 on ARO and SugarCrepe. These results suggest that image feature injection is more effective for improving compositionality than AdaIN, although the two methods are somewhat complementary. Through the improvements obtained using ViT-L/14 are a little bit different, both techniques improve performance over Variant #4. Combining both methods in Variant #7 leads to further improvements, with a more than 1% increase in average accuracy for ViT-B/32 and a 0.65% increase for ViT-L/14 over Variant #4. This highlights the importance of reducing unintended changes in synthetic images, as such changes could cause the model to rely on them for distinguishing positive and negative samples rather than the semantic differences specified in the corresponding captions. Examples of synthetic images in Figure A3 and A4 in the Appendix show how image feature injection helps mitigate these unintended changes.
@@
Effects of the margin loss and adaptive margin strategies. We compare three variants with distinct margin approaches: (1) Fixed (Variant #8): a fixed margin without adaptive adjustments; (2) Adaptive (Variant #9): the margin is calculated according to Eq. (4); and (3) Adaptive Inversed (Variant #10): a larger margin is applied to samples with higher similarity differences, representing an inverse version of Eq. (4) that prioritizes learning from easier samples. Comparing Variant #8 with Variant #7, we observe that #8 improves by over 1% for ViT-B/32 and 0.6% for ViT-L/14 in average accuracy, suggesting that margin loss effectively aids in compositional understanding. Between Variants #8, #9, and #10, we find that Variant #9 outperforms Variant #8 across all four benchmarks, with an average boost of over 0.3% for both ViT-B/32 and ViT-L/14. In contrast, Variant #10 results in a performance decline on three benchmarks and no improvement in average accuracy. These findings highlight the superior effectiveness of adaptive margins.
@@
Table 3: Performance (%) of SPARCL using different subsets of training samples. “Neg.” represents synthetic negative samples, “Pos.” represents synthetic positive samples, and “Prop.” represents the proportion of training samples.
Variant Neg. Pos. Prop.   ARO VL-CheckList SugarCrepe SugarCrepe++ Average
@@
Effects of synthetic positive samples. We compare the performance of training with only synthetic negative samples versus using both synthetic negative and positive samples in Table 3. Based on Variant #7, we train ViT-B/32 with these two combinations of training data. We observe nearly 3% decrease in average accuracy when using only synthetic negative samples, compared to using both synthetic negative and positive samples. This decrease is particularly noticeable on SugarCrepe++, ARO, and VL-CheckList. These results underscore the importance of incorporating synthetic positive samples in training, as they provide essential information about variations that maintain semantic consistency, which is critical for compositional understanding.
Influence of model size and training data size. To evaluate the impact of model size on compositional understanding, we compare the performance of ViT-B/32 and ViT-L/14 in Table 2. We find that ViT-L/14 only provides a modest improvement of about 0.5% of average accuracy over ViT-B/32. This suggests that increasing model size does not significantly enhance compositional understanding. To investigate the effect of training data size, we train ViT-B/32 using 20% and 50% of randomly sampled training data and report the results in Table 3. We observe that performance improves as the proportion of training data increases, indicating that a larger training set helps the model better capture compositional knowledge. However, the performance gain diminishes when the training data size increases from 50% to 100%. Perhaps generating or selecting high-quality data would be more effective than simplistic data scaling.
Additional experiments on hyperparameter analysis can be found in Appendix A3.
@@ 5 Conclusion
In this paper, we introduce SPARCL to enhance the compositional understanding capabilities of VLMs by generating and training with synthetic data. To tackle two key challenges in using synthetic data—namely, the difficulty of generating accurate variations and the inconsistency in cross-modal alignment quality—SPARCL integrates image feature injection into a T2I model to improve the quality of synthetic variations and introduces an adaptive margin loss to account for varying levels of cross-modal alignment for effectively learning nuanced distinctions. Experiments on four visual-language compositional understanding benchmarks demonstrate the effectiveness of SPARCL.
6 Acknowledgments
This work has been supported by the Nanyang Associate Professorship and the National Research Foundation Fellowship (NRF- NRFF13-2021-0006), Singapore. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the funding agencies.
@@ 6 Acknowledgments
  • Section A1 presents additional details about SPARCL.
@@
A1 Details of SPARCL
The prompts used as input to the LLM for generating negative and positive captions are presented in Figure A1.
@@
Data Synthesis. For caption generation, we utilize the Llama-2-Chat 13B model [^1] https://huggingface.co/meta-llama/Llama-2-13b-chat, with the temperature set to 0.9, top-k set to 100, and top-p set to 0.9 for sampling. For image generation, we use the LCM model [^2] https://huggingface.co/SimianLuo/LCM_Dreamshaper_v7 for its swift inference with few steps [61]. The pretrained CLIP ViT-L/14 [74] is used as the image feature extractor for injecting image features. We perform 8 inference steps with LCM to generate each image.
Hyperparameter Selection. First, we use only real training samples to select $\tau$ and $b$. The optimal values are determined by searching for the ones that minimize the training loss at the first training step, aiming to preserve the output distribution from the pretrained model. After searching, we set $\tau=0.01$ and $b=-30.0$. Next, we select the base learning rate, weight decay, and LoRA adapter rank based on performance on the COCO-2014 validation set, in which the model is trained exclusively on real samples. According to the performance on the validation set, these hyperparameter are set to a base learning rate of 0.01, weight decay of 0.5, and LoRA adapter rank of 16. Then, we construct a validation set composed of the CIFAR-10 [47] test set and a randomly selected 5% of samples from ARO-Attribute and ARO-Relation, to balance the performance on coarse-grained and fine-grained tasks. Using this validation set, we train the model on both real and synthetic samples and use the validation performance to determine the remaining hyperparameters: $m_{0}$, $\alpha$, $\beta$, $\gamma$ and $\lambda$. The effects of these hyperparameters are shown in Table A5.
@@
Table A1: Comparison of accuracy (%) between SPARCL and baselines on ARO and VL-CheckList. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Table A2: Comparison of accuracy (%) between SPARCL and baselines on SugarCrepe. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Table A3: Comparison of accuracy (%) between SPARCL and baselines on SugarCrepe++. “img” represents images, “cap” represents captions, “syn” represents synthetic data.
@@
Comparison with other images generation methods. We compare our image generation method with StyleAligned [27]. For a fair comparison, we use an ablated version #7 of SPARCL (Sec. 4.4, main paper) without the adaptive margin loss. Both methods use synthetic captions that we generate. As shown in Table A4, StyleAligned performs about 1% worse than our method on the four compositional benchmarks, which illustrates the effectiveness of image feature injection in SPARCL. In Figure A2, we show two synthetic images from StyleAligned, where it fails to align the generated content with the synthetic captions. We hypothesize that the diffusion trajectory of the real image imposes strong constraints on the image generation model, making StyleAligned difficult to edit the image to match the synthetic caption. This issue is similar to the zero-shot image editing methods [63, 6, 22], which provide incorrect guidance during model training and lead to limited improvements on compositional understanding tasks. Moreover, StyleAligned requires DDIM inversion to obtain the inverted diffusion trajectory from the real image, making it computationally expensive and impractical for large-scale image generation.
Table A4: Performance comparison (%) between SPARCL and StyleAligned.
Variant   ARO VL-CheckList SugarCrepe SugarCrepe++ Average
StyleAligned [27] 72.60 75.03 85.70 65.25 74.65
SPARCL (#7) 74.12 76.35 85.40 66.44 75.58
Figure A2: Examples of synthetic samples from StyleAligned. The algorithm did not alter the image content according to the caption.
@@
Effects of image feature injection. In Figure A3 and A4, we present examples of synthetic images to illustrate how image feature injection helps mitigate unintended changes. In Figure A3, we observe that feature injection helps to generate images with similar object size and viewing angle to the real image. For example, in (a), the real image depicts a wide shot of a girl, while the synthetic image without feature injection produces a close-up shot despite aligning with the caption. With feature injection, the synthetic image maintains a wide shot, resembling the real image. Similar effects are seen in (b) and (c). In (d), the synthetic image with feature injection preserves the viewing angle of the real image, whereas the one without feature injection deviates from it. In Figure A4, we observe that feature injection helps generate backgrounds that resemble the real image. For example, in (a), the real image and the synthetic image without feature injection depicts an outdoor street scene, creating a noticeable difference. With feature injection, the single-colored background makes the synthetic image more similar to the real one. In (b), the sky occupies much of background in the real image as well as the image generated with feature injection, whereas the one without feature injection shows little sky. Also, the basketball is present in both the real and the synthetic image with feature injection but not in the middle image. Similar effects are observed in (c) and (d). These examples show that image feature injection reduces unintended variations not captured by the caption, enhancing the usefulness of synthetic samples for training VLMs.
Figure A3: Examples of synthetic samples without and with image feature injection. In these examples, the image feature injection technique achieves alignment of the subject size and the viewing angle with those in real images.
Figure A4: Examples of synthetic samples without and with image feature injection. In these examples, the image feature injection primarily helps to generate backgrounds that resemble those in real images. For example, in (d), both the first and the third images show the ground, whereas the second image does not.
@@
Table A5: Performance of SPARCL with different hyperparameters. “ARO-Rel” refers to the ARO-Relation validation subset, and “ARO-Att” refers to the ARO-Attribute validation subset, both consisting of a randomly selected 5% of the full set, as described in Sec. A2.
@@
Effects of hyperparameters. Table A5 presents the performance of SPARCL with different hyperparameter settings. For $\lambda$, we observe that $\lambda=0.01$ achieves the highest average validation accuracy, leading us to select it for subsequent experiments. Similarly, for $\alpha$, the best performance is obtained with $\alpha=10.0$, which is used in other experiments. When evaluating different values of $m_{0}$, we find that $m_{0}=0.005$ yields the best results. Finally, we examine various combinations of $\beta$ and $\gamma$ and observe that $\beta=-0.02$ and $\gamma=1.0$ provide the best validation performance. Thus, this combination is selected as the optimal hyperparameter setting.